OpenStudy (anonymous):

what is the slope of the line that contains the points (0,4) and (4, -2)

OpenStudy (anonymous):

(x1,y1)=(0,4) and (x2,y2)=(4,-2) slope of the line =(y2-y1)/(x2-x1) hence=> -2-4/4-0 =>-6/4 =-3/2 :)

OpenStudy (anonymous):

thanks

OpenStudy (marihelenh):

\[\frac{ y _{2}-y _{2} }{ x _{2}-x _{1} }=\frac{ -2-4 }{ 4-0 }=\frac{ -6 }{ 4 }=\frac{ -3 }{ 2 }\]

OpenStudy (anonymous):

welcome @djsmart12